Air is being pumped into a spherical ballon at a rate of 10cm^3/min. Determine the rate at which the radius of the ballon is increasing when the radius of the ballon is 20cm
1
Expert's answer
2019-11-15T13:17:29-0500

The volume of the ball is calculated by the formula "V=\frac {4\pi r^{3}} {3}"

The rate of change of volume is "\dot{V}=4\pi r^2 \dot{r}"

Thus, "\dot{r}=\frac{\dot{V} }{4\pi r^2}=(10 cm^3/min)/(4\pi (20cm)^2)=" 0.00198943678864869169711104704216 cm/min